The ofice jet printer can copy Lisa's dissertation in 22 min. The laser jet printer can copy the same document in 10 min. if the two machines work together how long would they take to copy the dissertation?
Answer by ewatrrr(10682)

 
Hi,
Let x represent the time it would take the 2 machines working together
Question states***Per minute the Equalizer
1/22 + 1/10 = 1/x |multiplying thru by 22*10*x so as all denominators = 1
10x + 22x =10*22
32x = 220
x = 220/32
x = 6.875 min